I have a balloon that can hold 100 liters of air. If I blow up this balloon with 3 moles of oxygen gas at a pressure of 1 atmosphere, what is the temperature of the balloon?


1
Expert's answer
2021-05-17T04:08:40-0400

V = 100 L

n = 3 moles

p = 1 atm

Ideal gas law:

pV = nRT

R = 0.08206 L×atm/mol×K

"T = \frac{pV}{nR} \\

= \frac{1 \times 100}{3 \times 0.08206} = 406 \;K \\

T = 406 -273 = 133 \;ºC"

Answer: 133 °C